An electric heater is immersed in a well-stirred tank. Solvent at 19°C, which has a heat capacity of 2.1 kJ/(kg°C), is fed at a speed of 11 kg/h. The hot solvent is discharged at the same rate. The tank initially contains 180 kg of cold solvent at 10°C. The heat flow of the electric heater is 884 W. Do the following:

a) Draw the system with all its variables.
b) Write the differential equation along with its initial condition.
c) Calculate the time for the temperature of the solvent being discharged to be 63°C (write the answer in hours and with two decimals)
Answer #1

put T2= 63°6 -1697X10-5+ 63= 156.785 - 146 e † 93-785 :7146 e-1697x10-5 -1.697 10-5t е Hence 0-6423 = Hence t=26001.04 second